OpenStudy (anonymous):

David is selling floral arrangements. Each arrangement uses 1 vase and 12 roses. Each vase costs David $2.00. Let C be the total cost of the arrangement and r be the cost of 1 rose. Which equation should David use to find the total cost of each arrangement? C = 12r + 2 C = 2r + 12 12C = r + 12 12 = C + 2r

OpenStudy (mathstudent55):

If 1 rose costs r dollars, how much do 12 roses cost? What operation do you do to find that out?

OpenStudy (anonymous):

I dont know

OpenStudy (texaschic101):

example : if 1 orange costs $2, then 5 oranges cost 5(2) = $10 you multiply so if 1 rose costs r, then 12 roses cost ??

OpenStudy (mathstudent55):

1 rose costs r dollars 2 roses cost r + r = 2 * r = 2r dollars 3 roses cost r + r + r = 3 * r = 3r dollars 4 roses cost r + r + r + r = 4r dollars 5 roses cost 5r dollars 6 roses cost 6r dollars etc. 11 roses cost 11r dollars 12 roses cost ?

OpenStudy (texaschic101):

I can't explain it any easier then that ^^

OpenStudy (anonymous):

12 r dollars

OpenStudy (mathstudent55):

Correct. The 12 roses cost 12r dollars. That is only part of the total cost. The other part of the cost is the vase. The vase costs $2, and that cost of $2 has to be added to the 12r cost of the roses. Which expression in the choices shows that C, the cost, is $2 added to 12r?